OpenStudy (dawnr):

Can someone help me with this?! :// Sum of the equation(tan(2x+pi/4)=sqrt3 ) solutions for interval x€[0,pi) is?

OpenStudy (tkhunny):

Where is the tangent equal to sqrt(3)?

OpenStudy (dawnr):

i know that the arctan of sqrt is pi/3

OpenStudy (dawnr):

and that tan of pi/4 is 1 but idk what to do with it..

OpenStudy (triciaal):

I think 2x + pi/4 = 60

OpenStudy (dawnr):

how?

OpenStudy (triciaal):

(tan(2x+pi/4)=sqrt3 ) arctan of sqrt 3

OpenStudy (tkhunny):

No, 2x + pi/4 = pi/3 + kpi where k is any integer. There are LOTS of solutions.

OpenStudy (dawnr):

but i need to write the sum of the solutions in the interval above..

OpenStudy (tkhunny):

You must first find them all. 2x = pi/12 + kpi x = pi/24 + kpi/2 How many values of k put the corresponding value of x in the desired interval?

OpenStudy (dawnr):

0,pi/2,pi/3,pi/4? °_°

OpenStudy (dawnr):

also i never understood these : 2x = pi/12 + kpi x = pi/24 + kpi/2 are those always the same and how do i know which ones to use and when?

OpenStudy (triciaal):

 @tkhunny is this corrrect